OpenStudy (anonymous):

You can model the population of a certain city between the years 1955 and 1955 by the radical function: P(x)=^3√x-1940 Using this model, in what year was the population of that city 290,000? A. 1982 B. 1986 C. 1988 D. 1992

OpenStudy (anonymous):

Does the cube-root include 1940, or is it only over the x?

OpenStudy (anonymous):

Its included

OpenStudy (anonymous):

Okay, so you are given that you are looking for the population of the city equals 290,000. This means that P(x)=290,000 From here, you can solve for x.

OpenStudy (anonymous):

can you show me but don't give me the answer I want to see how it is done

OpenStudy (anonymous):

Sure. \[P(x)=\sqrt[3]{x-1940}\] \[P(x)=290,000\] Therefore, \[290,000=\sqrt[3]{x-1940}\]\[(290,000)^{3}=x-1940\]\[x=(209,000)^3+1940\]\[x=?\]

OpenStudy (anonymous):

Ok hold on I can do dis

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

um is the 209,000 supposed to be 290,000??

OpenStudy (anonymous):

Oh, yes, my apologies.

OpenStudy (anonymous):

its k

OpenStudy (anonymous):

would the answer be C??

OpenStudy (anonymous):

Yup. :)

OpenStudy (anonymous):

YAY I DID IT RIGHT XD

OpenStudy (anonymous):

Well you did the work but I still learned something :3

OpenStudy (anonymous):

I simply helped, and I am glad you learned something. :)

OpenStudy (anonymous):

Thanks again Jack ^v^

OpenStudy (anonymous):

No problem! :) Good luck with the rest of your homework!